OpenStudy (anonymous):

Sonji has $5.25 in nickels and dimes.The coin value can be modeled by the equation 0.10d + 0.05n = 5.25, where d represents the number of dimes and n represents the number of nickels. If Sonji has 21 dimes, how many nickels does she have? 35 42 63 105

OpenStudy (anonymous):

@10115658 Sorry for bugging you >.<

OpenStudy (anonymous):

replace d with 21 and solve for n

OpenStudy (anonymous):

n=63?

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Can you help with me with another? Lol

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A sequence is defined by the recursive function f(n + 1) = 1/3f(n). If f(3) = 9 , what is f(1) ?

OpenStudy (anonymous):